# RespoDiff: Dual-Module Bottleneck Transformation for Responsible & Faithful T2I Generation

### v1

@@ Abstract
The rapid advancement of diffusion models has enabled high-fidelity and semantically rich text-to-image generation; however, ensuring fairness and safety remains an open challenge. Existing methods typically improve fairness and safety at the expense of semantic fidelity and image quality. In this work, we propose RespoDiff, a novel framework for responsible text-to-image generation that incorporates a dual-module transformation on the intermediate bottleneck representations of diffusion models. Our approach introduces two distinct learnable modules: one focused on capturing and enforcing responsible concepts, such as fairness and safety, and the other dedicated to maintaining semantic alignment with neutral prompts. To facilitate the dual learning process, we introduce a novel score-matching objective that enables effective coordination between the modules. Our method outperforms state-of-the-art methods in responsible generation by ensuring semantic alignment while optimizing both objectives without compromising image fidelity. Our approach improves responsible and semantically coherent generation by ~20% across diverse, unseen prompts. Moreover, it integrates seamlessly into large-scale models like SDXL, enhancing fairness and safety. Code will be released upon acceptance.
@@ 1 Introduction
The rise of text-to-image diffusion models (T2I), such as Stable Diffusion, has transformed content creation and visual communication by generating high-quality visuals from text prompts (Podell et al., 2024; Rombach et al., 2022). However, these models risk reinforcing stereotypes or producing harmful content, which can lead to societal consequences (Luccioni et al., 2023; Perera and Patel, 2023; Rando et al., 2022; Schramowski et al., 2023). Ensuring a responsible workflow is critical to mitigating these risks.
Previous methods for responsible text-to-image (T2I) generation include prompt modification (Chuang et al., 2023; Ni et al., 2023), model fine-tuning (Gandikota et al., 2023; Shen et al., 2024), model editing (Gandikota et al., 2024), classifier guidance (Schramowski et al., 2023), and latent vector injection (Li et al., 2024). Despite advancements in responsible T2I generation, many existing approaches still compromise semantic fidelity and image quality, reducing their effectiveness in producing both responsible and faithful generation. To address these challenges, we propose RespoDiff, a novel framework for responsible T2I generation that introduces a dual-module transformation on the bottleneck representations of diffusion models. Specifically, given a responsible category, such as demographic attributes (e.g., gender) or safety factors – and its associated target concepts (e.g.,“man”, “woman” etc.), our approach learns independent transformations that steer the diffusion model toward target-aligned outputs while maintaining coherence with a neutral prompt (e.g., “a person”). These learned transformations can then be applied during inference to promote fairness and safety in T2I generation without compromising the underlying structure of the diffusion process. RespoDiff incorporates two learnable modules: ① Responsible Concept Alignment Module, which steers latent representations toward fair and safe outputs by learning transformations that align with responsible target concepts; and ② Semantic Alignment Module, which preserves consistency with neutral prompts, ensuring the generated images remain aligned with original prompts.
@@ 2 Related Work
T2I generation has transformed generative AI, enabling highly realistic image creation from text (Ho et al., 2020a; Ramesh et al., 2022; Rombach et al., 2022), but also raises ethical concerns, including the risk of generating harmful or inappropriate content (Cho et al., 2023; Luccioni et al., 2023).
Responsible Generation using Diffusion Models: In recent years, there has been a growing emphasis on methods to reduce biased and inappropriate content generation in diffusion models, such as Stable Diffusion. Several approaches focus on modifying input prompts by removing harmful or problematic terms (Ni et al., 2023; Schramowski et al., 2023), while others employ prompt-tuning techniques (Kim et al., 2023) or learn projection embeddings on prompt representations (Chuang et al., 2023) to filter out undesirable content. Some methods, such as those by Gandikota et al. (2023), Huang et al. (2023) and Zhang et al. (2024) attempt to erase unsafe concept representations from the diffusion models, though these techniques may negatively affect the model’s original performance. Similarly, Shen et al. (2024) addresses biases by fine-tuning specific parts of the model weights, but such approaches require additional training for each prompt or domain. In contrast, our method avoids prompt-specific fine-tuning and generalizes effectively across diverse and unseen prompts, including profession-specific ones. Alternative strategies such as those by Friedrich et al. (2023), Parihar et al. (2024), and Schramowski et al. (2023) use classifier-free guidance to steer the generation process away from undesirable content without requiring extra training. While some methods propose efficient closed-form solutions for embedding matrices to ensure responsible content generation (Chuang et al., 2023; Gandikota et al., 2024), they lack the adaptability and fine-grained control over image generation offered by our approach.
@@ 3 Preliminaries
In this section, we provide the necessary background regarding diffusion models and the scoring functions which forms the foundation of our model design.
Diffusion Models: Diffusion models (Ho et al., 2020b; Sohl-Dickstein et al., 2015) are likelihood-based generative models inspired by nonequilibrium thermodynamics (Song and Ermon, 2019). The model learns a denoising process that transforms random noise into samples from original data distribution, $p_{\text{data}}$. The process involves gradually corrupting training data with Gaussian noise in a forward process, where an initial sample $\bm{x}_{0}\sim p_{\text{data}}$ is progressively noised into $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ through a Markovian process as follows :
@@
where $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ is parameterized using a noise prediction network $\epsilon_{\theta}(\bm{x}_{t},t)$. After training, generation in diffusion models involves sampling from $p_{\bm{\theta}}(\bm{x}_{0})$, starting with a noise sample $\bm{x}_{T}\sim p(\bm{x}_{T})$ and recovering $\bm{x}_{0}\sim p_{\text{data}}$ using an SDE/ODE solver (e.g., DDIM (Song et al., 2021)). These models learn the transition probabilities $p(\bm{x}_{t-1}|\bm{x}_{t})$, defined as follows:
@@
Diffusion Scoring Function: The noise prediction network $\epsilon_{\theta}(\bm{x}_{t},t)$ iteratively estimates the noise $\epsilon$ to generate $\bm{x}_{0}$ from $\bm{x}_{T}$ and approximates the score function (Ho et al., 2020b; Song et al., 2020), given by $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\approx-\epsilon_{\theta}(\bm{x}_{t},t)/\sigma_{t}$, where $\sigma_{t}$ is the noise level at time step $t$ and $p_{t}$ is the marginal distribution of the samples noised to time $t$. Following the score function direction guides samples back to the data distribution.
@@ 4.2 Responsible Concept Alignment Module
It is important to note that during this stage, the loss computation considers only $f_{\text{resp}}$, and not $\hat{f}$, as the sole objective is to steer the RAM towards aligning with the target concept. Following Song et al. (2020), the objective in Eq.˜4 can be viewed as minimizing the gradient disparity between the neutral and target diffusion trajectories, effectively guiding the latent space toward the target concept at each timestep. As a result, optimizing this score-matching loss enables the RAM to transform the latent representation such that the generated images are effectively aligned with the desired target concept. Unlike prior methods that rely on implicit reconstructions using pre-existing target images, our loss function explicitly supervises the transformation in the model’s latent space by precisely aligning the diffusion trajectories of neutral and target concepts. By leveraging the neutral denoised latent as a stable reference point, our approach ensures that the modification is structurally guided rather than inferred indirectly from external data.
@@ 5.1 Fair Generation
Evaluation setting: We evaluate our method on the Winobias benchmark (Zhao et al., 2018), following the approaches in (Gandikota et al., 2024; Li et al., 2024; Orgad et al., 2023), which includes 36 professions with known gender biases. We learn transformations as constant functions linearly added to bottleneck activations, optimizing over 5000 iterations with batch size 1, as in Section˜4.2. Unlike Gandikota et al. (2024) and Shen et al. (2024), our approach does not rely on profession-specific data during training. Instead, we utilize the prompt “a person” to learn generalized directions that are applicable across professions. For fair comparison, we adopt the experimental setup from Li et al. (2024) to evaluate fairness. Five prompts per profession are used, including templates like “A photo of a $\langle\text{profession}\rangle$”. We also extend our evaluation to the Gender+ and Race+ Winobias datasets (Li et al., 2024), which introduce terms like “successful” to trigger stereotypical biases (Gandikota et al., 2024). Additional dataset details are in Section˜7.4.1 of Appendix.
Metrics: We perform quantitative and qualitative analysis to evaluate the performance of our proposed approach. We employ the modified average deviation ratio (DevRat), as defined in Li et al. (2024) to quantify the fairness of the generated images. We also measure the alignment of the generated images with the Winobias prompts (WinoAlign). Additionally, we assess image fidelity using the FID score (Heusel et al., 2017) on the COCO-30K validation set (FID (30K)), while image-text alignment is measured with the CLIP score (Radford et al., 2021a) using COCO-30K prompts under fair transformations (CLIP (30K)). Further details are provided in Section˜7.4.6 of Appendix.
@@
Results: We compare our approach against baselines including Stable Diffusion (SD) (Rombach et al., 2022), FDF (Shen et al., 2024), SDisc (Li et al., 2024) and BAct (Parihar et al., 2024) with additional details in Section˜7.4.2. Tables˜2 and 2 present a comparison of our approach to various baseline methods, focusing on various metrics across both gender and race biases. We further analyze extended biases within these categories, with additional comparisons provided in Section˜7.4.7. RespoDiff consistently achieves the lowest average deviation ratio in both gender and race biases, even in challenging settings, highlighting its superior performance in mitigating biases across different professions. As observed in Section˜7.4.8 of the Appendix, our method effectively eliminates gender and racial biases in a range of professions compared to Stable Diffusion. Although FDF performs better in certain professions like Secretary, likely due to training on profession-specific images, our approach improves fairness across all professions on average without being explicitly trained on profession-specific prompts. This highlights our model’s strong generalization ability across different professions. Additionally, RespoDiff remains robust to bias-inducing prompts with phrases such as ‘successful’, as detailed in Section˜7.4.7. We also provide additional comparisons in Section˜7.4.3.
@@
Qualitative analyses in Fig.˜2 further support our quantitative findings. RespoDiff effectively modifies Stable Diffusion outputs to align with sensitive subconcepts like “woman”, “black race” etc. using learned transformations while preserving profession-specific attributes. Additional results are available in Section˜7.8. We also present additional qualitative comparisons of our approach with hard prompting and SDisc in Section˜7.4.4 and Section˜7.4.5, respectively.
@@ 5.2 Safe Generation
Metrics: To assess inappropriateness, we utilize a combination of predictions from the Q16 classifier and the NudeNet classifier on the generated images, in line with the approaches presented in Gandikota et al. (2023); Schramowski et al. (2023); Li et al. (2024). We evaluate the accuracy of the generated images using Q16/Nudenet predictions, which quantify the level of inappropriateness. We also compute the FID and CLIP scores to assess image fidelity and image-text alignment using the COCO-30K prompts. We compare our approach against baselines such as SD, SDisc, ESD (Gandikota et al., 2023) and SLD (Schramowski et al., 2023). Further details are provided in Section˜7.5.2.
@@
SD Rombach et al. (2022) 0.27 14.09 31.33
SDisc Li et al. (2024) 0.27 15.98 31.03
SLD Schramowski et al. (2023) 0.20 18.76 29.75
ESD Gandikota et al. (2023) 0.32 13.68 30.43
@@
Despite training on a limited safety set $\mathcal{S}_{\text{safe}}=\{\text{violence},\text{nudity}\}$, RespoDiff generalizes well to other safety-critical categories, as detailed in Section˜7.5.3. Qualitative results are shown in Fig.˜3. We further assess FID and CLIP scores, demonstrating that our approach maintains image quality comparable to Stable Diffusion on COCO-30K while achieving superior image-text alignment. Notably, our method not only enhances safety beyond SLD, which is designed specifically for safety, but also outperforms it in image-text alignment. While ESD and SDisc achieve higher image quality, our approach effectively filters inappropriate content without significant degradation in visual fidelity. Additional results are provided in Section˜7.8.
@@ 5.3 Responsible Generation using SDXL
In this section, we explore the applicability of RespoDiff to large-scale generative models, specifically SDXL, for responsible T2I generation. We train our method using SDXL for the fair generation task and evaluate its effectiveness by comparing the results against the baseline SDXL model. The quantitative results are reported in Table˜4 and Table˜5.
As shown in Table˜4 and Table˜5, RespoDiff significantly mitigates biases related to gender and race that are present in SDXL, achieving substantial improvements. Importantly, our method preserves both the visual quality and image-text alignment of the original SDXL model. We also provide qualitative results demonstrating the transformation of concepts such as woman, Black race, Asian race, and White race in Fig.˜4 and Section˜7.8 of the Appendix. Notably, RespoDiff accurately steer generated images toward target concepts while maintaining consistency with the original content.

### v2

@@ Abstract
The rapid advancement of diffusion models has enabled high-fidelity and semantically rich text-to-image generation; however, ensuring fairness and safety remains an open challenge. Existing methods typically improve fairness and safety at the expense of semantic fidelity and image quality. In this work, we propose RespoDiff, a novel framework for responsible text-to-image generation that incorporates a dual-module transformation on the intermediate bottleneck representations of diffusion models. Our approach introduces two distinct learnable modules: one focused on capturing and enforcing responsible concepts, such as fairness and safety, and the other dedicated to maintaining semantic alignment with neutral prompts. To facilitate the dual learning process, we introduce a novel score-matching objective that enables effective coordination between the modules. Our method outperforms state-of-the-art methods in responsible generation by ensuring semantic alignment while optimizing both objectives without compromising image fidelity. Our approach improves responsible and semantically coherent generation by ~20% across diverse, unseen prompts. Moreover, it integrates seamlessly into large-scale models like SDXL, enhancing fairness and safety .The project page is available at https://vssilpa.github.io/respodiff_project_page.
@@ 1 Introduction
Models such as Stable Diffusion v1.4 (SDv1.4), SDXL, FLUX, and SD3 exemplify recent advancements in text-to-image (T2I) generation, revolutionizing content creation and visual communication by generating high-quality visuals from text prompts (Podell et al., 2024; Rombach et al., 2022). However, these models risk reinforcing stereotypes or producing harmful content, which can lead to societal consequences (Luccioni et al., 2023; Perera and Patel, 2023; Rando et al., 2022; Schramowski et al., 2023). Ensuring a responsible workflow is critical to mitigating these risks.
Previous methods for responsible text-to-image (T2I) generation include prompt modification (Chuang et al., 2023; Ni et al., 2024), model fine-tuning (Gandikota et al., 2023; Shen et al., 2024), model editing (Gandikota et al., 2024), classifier guidance (Schramowski et al., 2023), and latent vector injection (Li et al., 2024). Despite advancements in responsible T2I generation, many existing approaches still compromise semantic fidelity and image quality, reducing their effectiveness in producing both responsible and faithful generation. To address these challenges, we propose RespoDiff, a novel framework for responsible T2I generation that introduces a dual-module transformation on the bottleneck representations of diffusion models. Specifically, given a responsible category, such as demographic attributes (e.g., gender) or safety factors – and its associated target concepts (e.g.,“man”, “woman” etc.), our approach learns independent transformations that steer the diffusion model toward target-aligned outputs while maintaining coherence with a neutral prompt (e.g., “a person”). These learned transformations can then be applied during inference to promote fairness and safety in T2I generation without compromising the underlying structure of the diffusion process. RespoDiff incorporates two learnable modules: ① Responsible Concept Alignment Module, which steers latent representations toward fair and safe outputs by learning transformations that align with responsible target concepts; and ② Semantic Alignment Module, which preserves consistency with neutral prompts, ensuring the generated images remain aligned with original prompts.
@@ 2 Related Work
T2I generation has transformed generative AI, enabling highly realistic image creation from text (Ho et al., 2020; Ramesh et al., 2022; Rombach et al., 2022), but also raises ethical concerns, including the risk of generating harmful or inappropriate content (Cho et al., 2023; Luccioni et al., 2023).
Responsible Generation using Diffusion Models: In recent years, there has been a growing emphasis on methods to reduce biased and inappropriate content generation in diffusion models, such as Stable Diffusion. Several approaches focus on modifying input prompts by removing harmful or problematic terms (Ni et al., 2024; Schramowski et al., 2023), while others employ prompt-tuning techniques (Kim et al., 2023) or learn projection embeddings on prompt representations (Chuang et al., 2023) to filter out undesirable content. Some methods, such as those by Gandikota et al. (2023), Huang et al. (2024) and Zhang et al. (2024) attempt to erase unsafe concept representations from the diffusion models, though these techniques may negatively affect the model’s original performance. Similarly, Shen et al. (2024) addresses biases by fine-tuning specific parts of the model weights, but such approaches require additional training for each prompt or domain. In contrast, our method avoids prompt-specific fine-tuning and generalizes effectively across diverse and unseen prompts, including profession-specific ones. Alternative strategies such as those by Friedrich et al. (2023), Parihar et al. (2024), and Schramowski et al. (2023) use classifier-free guidance to steer the generation process away from undesirable content without requiring extra training. While some methods propose efficient closed-form solutions for embedding matrices to ensure responsible content generation (Chuang et al., 2023; Gandikota et al., 2024), they lack the adaptability and fine-grained control over image generation offered by our approach.
@@ 3 Preliminaries
In this section, we provide the necessary background regarding diffusion models and the scoring functions which form the foundation of our model design.
Diffusion Models: Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) are likelihood-based generative models inspired by nonequilibrium thermodynamics (Song and Ermon, 2019). The model learns a denoising process that transforms random noise into samples from original data distribution, $p_{\text{data}}$. The process involves gradually corrupting training data with Gaussian noise in a forward process, where an initial sample $\bm{x}_{0}\sim p_{\text{data}}$ is progressively noised into $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ through a Markovian process as follows :
@@
where $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ is parameterized using a noise prediction network $\epsilon_{\theta}(\bm{x}_{t},t)$. After training, generation in diffusion models involves sampling from $p_{\bm{\theta}}(\bm{x}_{0})$, starting with a noise sample $\bm{x}_{T}\sim p(\bm{x}_{T})$ and recovering $\bm{x}_{0}\sim p_{\text{data}}$ using an SDE/ODE solver (e.g., DDIM (Song et al., 2021a)). These models learn the transition probabilities $p(\bm{x}_{t-1}|\bm{x}_{t})$, defined as follows:
@@
Diffusion Scoring Function: The noise prediction network $\epsilon_{\theta}(\bm{x}_{t},t)$ iteratively estimates the noise $\epsilon$ to generate $\bm{x}_{0}$ from $\bm{x}_{T}$ and approximates the score function (Ho et al., 2020; Song et al., 2021b), given by $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\approx-\epsilon_{\theta}(\bm{x}_{t},t)/\sigma_{t}$, where $\sigma_{t}$ is the noise level at time step $t$ and $p_{t}$ is the marginal distribution of the samples noised to time $t$. Following the score function direction guides samples back to the data distribution.
@@ 4.2 Responsible Concept Alignment Module
It is important to note that during this stage, the loss computation considers only $f_{\text{resp}}$, and not $\hat{f}$, as the sole objective is to steer the RAM towards aligning with the target concept. Following Song et al. (2021b), the objective in Eq.˜4 can be viewed as minimizing the gradient disparity between the neutral and target diffusion trajectories, effectively guiding the latent space toward the target concept at each timestep. As a result, optimizing this score-matching loss enables the RAM to transform the latent representation such that the generated images are effectively aligned with the desired target concept. Unlike prior methods that rely on implicit reconstructions using pre-existing target images, our loss function explicitly supervises the transformation in the model’s latent space by precisely aligning the diffusion trajectories of neutral and target concepts. By leveraging the neutral denoised latent as a stable reference point, our approach ensures that the modification is structurally guided rather than inferred indirectly from external data.
@@ 5.1 Fair Generation
Evaluation Setting: We evaluate our method on the Winobias benchmark (Zhao et al., 2018), following the approaches in (Gandikota et al., 2024; Li et al., 2024; Orgad et al., 2023), which includes 36 professions with known gender biases. We learn transformations as constant functions linearly added to bottleneck activations, optimizing over 5000 iterations with batch size 1, as in Section˜4.2. Unlike Gandikota et al. (2024) and Shen et al. (2024), our approach does not rely on profession-specific data during training. Instead, we utilize the prompt “a person” to learn generalized directions that are applicable across professions. For fair comparison, we adopt the experimental setup from Li et al. (2024) to evaluate fairness. Five prompts per profession are used, including templates like “A photo of a $\langle\text{profession}\rangle$”. We also extend our evaluation to the Gender+ and Race+ Winobias datasets (Li et al., 2024), which introduce terms like “successful” to trigger stereotypical biases (Gandikota et al., 2024). Additional dataset details are in Section˜8.4.1 of Appendix.
Metrics: We perform quantitative and qualitative analysis to evaluate the performance of our proposed approach. We employ the modified average deviation ratio (DevRat), as defined in Li et al. (2024) to quantify the fairness of the generated images. We also measure the alignment of the generated images with the Winobias prompts (WinoAlign). Additionally, we assess image fidelity using the FID score (Heusel et al., 2017) on the COCO-30K validation set (FID (30K)), while image-text alignment is measured with the CLIP score (Radford et al., 2021) using COCO-30K prompts under fair transformations (CLIP (30K)). Further details are provided in Section˜8.4.6 of Appendix.
@@
Results: We compare our approach against baselines including Stable Diffusion (SD) (Rombach et al., 2022), FDF (Shen et al., 2024), SDisc (Li et al., 2024) and BAct (Parihar et al., 2024) with additional details in Section˜8.4.2. Tables˜2 and 2 present a comparison of our approach to various baseline methods, focusing on various metrics across both gender and race biases. We further analyze extended biases within these categories, with additional comparisons provided in Section˜8.4.7. RespoDiff consistently achieves the lowest average deviation ratio in both gender and race biases, even in challenging settings, highlighting its superior performance in mitigating biases across different professions. As observed in Section˜8.4.8 of the Appendix, our method effectively eliminates gender and racial biases in a range of professions compared to Stable Diffusion. Although FDF performs better in certain professions like Secretary, likely due to training on profession-specific images, our approach improves fairness across all professions on average without being explicitly trained on profession-specific prompts. This highlights our model’s strong generalization ability across different professions. Additionally, RespoDiff remains robust to bias-inducing prompts with phrases such as ‘successful’, as detailed in Section˜8.4.7. We also provide additional comparisons in Section˜8.4.3.
@@
Qualitative analyses in Fig.˜2 further support our quantitative findings. RespoDiff effectively modifies Stable Diffusion outputs to align with sensitive subconcepts like “woman”, “black race” etc. using learned transformations while preserving profession-specific attributes. Additional results are available in Section˜8.9. We also present additional qualitative comparisons of our approach with hard prompting and SDisc in Section˜8.4.4 and Section˜8.4.5, respectively.
@@ 5.2 Safe Generation
Metrics: To assess inappropriateness, we utilize a combination of predictions from the Q16 classifier and the NudeNet classifier on the generated images, in line with the approaches presented in Gandikota et al. (2023); Schramowski et al. (2023); Li et al. (2024). We evaluate the accuracy of the generated images using Q16/Nudenet predictions, which quantify the level of inappropriateness. We also compute the FID and CLIP scores to assess image fidelity and image-text alignment using the COCO-30K prompts. We compare our approach against baselines such as SD, SDisc, ESD (Gandikota et al., 2023) and SLD (Schramowski et al., 2023). Further details are provided in Section˜8.5.2.
@@
SD (Rombach et al., 2022) 0.27 14.09 31.33
SDisc (Li et al., 2024) 0.27 15.98 31.03
SLD (Schramowski et al., 2023) 0.20 18.76 29.75
ESD (Gandikota et al., 2023) 0.32 13.68 30.43
@@
Despite training on a limited safety set $\mathcal{S}_{\text{safe}}=\{\text{violence},\text{nudity}\}$, RespoDiff generalizes well to other safety-critical categories, as detailed in Section˜8.5.3. Qualitative results are shown in Fig.˜3. We further assess FID and CLIP scores, demonstrating that our approach maintains image quality comparable to Stable Diffusion on COCO-30K while achieving superior image-text alignment. Notably, our method not only enhances safety beyond SLD, which is designed specifically for safety, but also outperforms it in image-text alignment. While ESD and SDisc achieve higher image quality, our approach effectively filters inappropriate content without significant degradation in visual fidelity. Additional results are provided in Section˜8.9.
@@ 5.3 Responsible Generation using SDXL
In this section, we explore the applicability of RespoDiff to large-scale generative models, specifically SDXL, for responsible T2I generation. We train our method using SDXL for the fair generation task and evaluate its effectiveness by comparing the results against the baseline SDXL model. The quantitative results are reported in Table˜4 and Table˜5. We also train RespoDiff using SDXL for safe generation and the results are provided in Section˜8.5.4.
As shown in Table˜4 and Table˜5, RespoDiff significantly mitigates biases related to gender and race that are present in SDXL, achieving substantial improvements. Importantly, our method preserves both the visual quality and image-text alignment of the original SDXL model. We also provide qualitative results demonstrating the transformation of concepts such as woman, Black race, Asian race, and White race in Fig.˜4 and Section˜8.9 of the Appendix. Notably, RespoDiff accurately steer generated images toward target concepts while maintaining consistency with the original content.